OpenStudy (anonymous):

what steps would I take to solve the problem: Solve each equation on the interval 0

OpenStudy (anonymous):

Do you mean \(0<x<2\pi\) or do you mean \(0 < \alpha < 2\pi\)?

OpenStudy (anonymous):

the 2nd one

OpenStudy (anonymous):

But there are no alphas in your problems...

OpenStudy (anonymous):

that why im confused as well lol maybe it is the first one

OpenStudy (anonymous):

Let's just assume that it is the first one, because that would make more sense lol

OpenStudy (anonymous):

So, for the first one.\[2\sin(x)+\sqrt{2}=0\]\[2\sin(x)=-\sqrt{2}\]\[\sin(x)=-\frac{\sqrt{2}}{2}\]Now, use your trig tables to find what values of x satisfy that while making sure \(0 < x < 2\pi\). Get it?

OpenStudy (anonymous):

oh so you pretty much want to get the sin alone

OpenStudy (anonymous):

what if there is sin and cos in the problem. ex:8-12sin^2x=4cos^2x

OpenStudy (anonymous):

Yeah. When solving trig equations, you want to (1) get all the trig stuff into 1 trigonometric function. So, if you have both a sine and a cosine or something, you want to (usually) get it all in terms of sine or cosine. (2) You want to solve for the trig function.

OpenStudy (anonymous):

For that kind of thing, that's where you have to use your identities. For instance, in that question you just posted, trying using your pythagorean identities. \[\implies8 - 12\sin^2(x)=4\cos^2(x)\]\[\implies8-12\sin^2(x)=4(1-\sin^2(x))\]\[\implies0=8\sin^2(x)-4\]\[\implies\frac12=\sin^2(x)\]Does this make sense?

OpenStudy (anonymous):

ahh yes I finally understand :D thanks so much!

OpenStudy (anonymous):

You're welcome. :)